OpenStudy (anonymous):

If I have a function: y=K*e^x is there a way to rewrite it as y=C*A^x where A does not equal e?

OpenStudy (experimentx):

let, A^i = e i = 1/lnA

OpenStudy (experimentx):

it will be the same nevertheless, however I am not sure about K and C

OpenStudy (anonymous):

I´m just wondering if e is such a magical number to appear here and there or if its artificial and we could use a different set of numbers to express the same thing/physical law.

OpenStudy (experimentx):

i think e is quite an special number <-- especially when rate of change depends on initial value. also it has some special property (base of natural log), i think it is best to leave things with e's

OpenStudy (experimentx):

also stretching or compressing exponential function gives e at some point

OpenStudy (anonymous):

Hmmm what is so special about e? I only know that the derivative of e^x is again e^x which is quite nice and that I can rewrite e^(ix) in terms of sine and cosine but is there even more to e?

OpenStudy (experimentx):

LOL ... not really sure!!

OpenStudy (anonymous):

xD I thought there might be some additional stuff I might not know about.^^

OpenStudy (beginnersmind):

Medals 0 You can do it by writing e as a^(1/ln(a)) and apply exponential identities but why would you want to? e is such a nice number :)

OpenStudy (experimentx):

e's the best ... LOL

OpenStudy (beginnersmind):

From a practical point of view, if you need to differentiate or integrate your function down the road, would you rather deal with e^x or A^x?

OpenStudy (anonymous):

@beginnersmind I was wondering if the appearance of e in some laws of physics was because somewhat had the hearts for e or because there is no other way to state that law. Might have been just some physicist who loved e and we could rewrite some laws. Ok than I guess e really is that great. :)

OpenStudy (experimentx):

comes naturally from \( \int 1/x dx \)

OpenStudy (anonymous):

f I integrate or differentiate I am always happy to encounter e ;) But if I just add and multiply I could live without e ;) So it depends on the field.

OpenStudy (beginnersmind):

well, a lot of laws are actually of the form e^Cx, so I'm not sure e is especially preferred by nature in those cases. It seems to be notational .

OpenStudy (anonymous):

So I could easily rewrite y=K*e^(Cx) with some other base?

OpenStudy (experimentx):

i've usually encountered in decay equation and distribution function. decay equation <--- dN/dt = N <-- depends on initial value distribution function --> didn't understand

OpenStudy (anonymous):

I started wondering while looking at the decay equation like 10 Minutes ago.^^

OpenStudy (beginnersmind):

@experimentX in the decay function the choice of the base is somewhat arbitrary. Say you have f(t)=e^(-Ct). You could just as easily write f(t)-2^(-Kt), with a different constant.

OpenStudy (experimentx):

why ... we choose it as natural base for log while integrating <--- must be some reason.

OpenStudy (beginnersmind):

I'm saying e^(-Ct) and 2^(Ct/ln2) are the same function. They take the same values.

OpenStudy (beginnersmind):

2^(-Ct/ln2) that is

OpenStudy (anonymous):

But If you rewrite it using the ln than e is still in there (in that function). So you can rewrite in a way that you can not see e but it is hidden in the ln.

OpenStudy (experimentx):

we could use the same logic everywhere, the point is why e so that there is no logs on the power?? 1/ ln 2 ??

OpenStudy (beginnersmind):

Well, there's a constant, which is an experimentally determined number. When you use e as the base the constant is C. When you use 2 it's K=C/ln2. If you actually measure half-life you're measuring K, so ln2 isn't really "hidden" there.

OpenStudy (anonymous):

Ops right. ln2 is just some number - there is not an e hidden. So we could really rewrite equations that fit that pattern.

OpenStudy (experimentx):

not really sure if i am understanding ..:(

OpenStudy (anonymous):

@experimentX Which part? You wrote the same thing as beginnersmind with the rewriting 1/lnA or now 1/ln2.

OpenStudy (experimentx):

e^x/nothing <--- why nothing in this case?? must be some special property of e A^x/lnA

OpenStudy (anonymous):

Why e^x/nothing? Who wrote that where?

OpenStudy (experimentx):

usually equation comes that way when we integrate 1/x dx Oo, i think i need to review decay equation, since i ignored \( \lamda \) factor completely.

OpenStudy (experimentx):

LOL ... seems like only e is nice to deal with,

OpenStudy (anonymous):

:) I guess than everything is alright. What a nice and interesting discussion. If you follow mathematics and physics in class it seems that e is mysteriously everywhere but apparently it is that way because some people are secretly working for e and we could write it in a different way. :) Nevertheless e is a great number for calculus.

OpenStudy (experimentx):

yeah ... that i agree!! makes nice, easy and clean.

OpenStudy (beginnersmind):

To be fair, there are situations where e appears in its own right. E.g. you can't rewrite e^i*pi=-1 with any number. (I think)